OpenStudy (swissgirl):

Prove by the principal of mathematical Induction: \( \frac{1} {2!} + \frac{2} {3!} +...+ \frac{n} {(n+1)!} = 1 - \frac{1} {(n+1)!} \)

OpenStudy (anonymous):

If you type in "huge" or "large" in the equation editor, it'll make all that bigger fwiw. :D So how would you start off?

OpenStudy (swissgirl):

The first step wld be to verify that this is true witthe lowest value of n which is 1 in thsi case

OpenStudy (anonymous):

Right, our Basis case right?

OpenStudy (swissgirl):

yup

OpenStudy (swissgirl):

which obv is true

OpenStudy (anonymous):

Next we need to start our IH (induction Hypothesis). One thing that I always do is the following: Let P(n) be the statement "insert huge statement here"

OpenStudy (swissgirl):

then P(n+1) is true

OpenStudy (anonymous):

Then i follow in the IH like so: Spse P(n) is try, show that P(n+1) is true for any arbitrary n.

OpenStudy (swissgirl):

ya so that is my difficulty

OpenStudy (anonymous):

Since the LHS is in a series, when we add "one" we add a whole new term. therefore we would add an additional: \[\huge + {n+1 \over (n+2)!} = 1 - {1 \over (n+1)!}\]

OpenStudy (anonymous):

so on my LHS just imagine that there is everything you wrote from the LHS before my "plus" sign. Does that make sense?

OpenStudy (swissgirl):

welll y didnt u add a +1 on the right hand side?

OpenStudy (anonymous):

Oh whoops... you're right. Good catch.

OpenStudy (anonymous):

But since we're suposing that the Basis was true, P(n), we can replace the whole series representation with: \[\huge 1-{1 \over (n+1)!}\] and can then combine everything together now (correctly) \[\huge 1 - {1 \over (n+1)!} + {n+1 \over (n+2)!} = 1 - {1 \over (n+2)!}\]

OpenStudy (anonymous):

Think you can take it from there?

OpenStudy (swissgirl):

idk i didnt follow at all

OpenStudy (callisto):

Let p(n) be the statement \(\frac{1}{2!} + \frac{2}{3!} + ... + \frac{n}{(n+1)!} = 1-\frac{1}{(n+1)!}\) Consider p(1) LS = \( \frac{1}{(1+1)!} = \frac{1}{2}\) RS = \(1-\frac{1}{(1+1)!} = \frac{1}{2}\) So, P(1) is true Suppose p(k) is true for some positive integers k. LS of P(k+1) = \(\frac{1}{2!} + \frac{2}{3!} + ... + \frac{k}{(k+1)!} +\frac{k+1}{[(k+1)+1]!}\) = \( 1-\frac{1}{(k+1)!}+\frac{k+1}{(k+2)!}\) = \( 1-\frac{(k+2) - (k+1)}{(k+2)!}\) = \(1 -\frac{1}{(k+2)!}\) = RS of P(k+1) I think you can continue from here?

OpenStudy (swissgirl):

I didnt follow one point. The third line where you everything under (k+2)! How did u get (k+2)-(K+1). I just suck at sequences like I never really learnt it

OpenStudy (callisto):

(k+2)! = (k+2)(k+1)! \(1-\frac{(1}{(k+1)!}+\frac{(k+1)}{(k+2)!} \) \(=1-\frac{1(k+2)}{(k+2)(k+1)!}+\frac{(k+1)}{(k+2)!}\) \(=1-\frac{(k+2)}{(k+2)!}+\frac{(k+1)}{(k+2)!}\) \(=1-\frac{(k+2)-(k+1)}{(k+2)!}\)

OpenStudy (swissgirl):

ohhhhhhh i seee lol

OpenStudy (swissgirl):

hahahahaha

OpenStudy (swissgirl):

Ok the rest of the proof was really clear. THANKKKKKSSSSS

OpenStudy (swissgirl):

THAT WAS SERIOUSLY AWESOME

OpenStudy (swissgirl):

THANKS :))))))))))))))))

OpenStudy (callisto):

Welcome, and thanks! :)